OpenStudy (anonymous):

Evaluate the arcos Sqrt (3)/2?

OpenStudy (anonymous):

\[\cos ^{-1}\left(\frac{\sqrt{3}}{2}\right)=\frac{\pi }{6}=30 {}^{\circ} \]

OpenStudy (anonymous):

Thank You!

OpenStudy (jhannybean):

Draw out a triangle for yourself.

OpenStudy (jhannybean):